OpenStudy (nicholausblackmon):

Can someone help me figure out how to solve this question? In 4 + In (3x)=2

OpenStudy (nicholausblackmon):

Note: I have absolutely no idea how to. I completely forgot

OpenStudy (solomonzelman):

You need to know: \(\color{#000000 }{ \displaystyle \ln(a)+\ln(b)=\ln(a\cdot b) }\) \(\color{#000000 }{ \displaystyle \ln(z) =y\quad \Longrightarrow \quad z=e^y }\)

OpenStudy (solomonzelman):

Use the first rule to combine the logarithms on the left side.

OpenStudy (nicholausblackmon):

can you explain how i should plug them in because like i said, I have absolutely no idea. This was given to me as bonus for my teacher last week and she said that i could still do it

OpenStudy (solomonzelman):

In your case, a is 4 b is 3x

OpenStudy (solomonzelman):

If you know; \(\color{#000000 }{ \displaystyle \ln(a)+\ln(b)=\ln(a\cdot b) }\) Then, \(\color{#000000 }{ \displaystyle \ln(4)+\ln(3x)=? }\)

OpenStudy (nicholausblackmon):

so its in (4) + in(3x)=2?

OpenStudy (solomonzelman):

"in" ?

OpenStudy (solomonzelman):

That is ell, not i. "ln" denotes/abbreviates - natural logarithm

OpenStudy (nicholausblackmon):

ln. sorry

OpenStudy (solomonzelman):

Can you apply the [first] rule that I provided?

OpenStudy (nicholausblackmon):

Im on my brothers account btw. Im in 7th grade and have absolutely no idea what ln means. Can you please explain it to me?

OpenStudy (solomonzelman):

have you heard of a function \(e^x\) ?

OpenStudy (nicholausblackmon):

yes, i have

OpenStudy (solomonzelman):

Ok, very good

OpenStudy (solomonzelman):